OpenStudy (anonymous):

cos^2x(1+tan^2x)=1 is this an identity? explain why

OpenStudy (psymon):

Well, do you know this identity? \[\sin ^{2}x+\cos ^{2}x=1\]

OpenStudy (unklerhaukus):

tan=sin/cos tan^2=sin^2/cos^2